### v2

@@ 3.2 The GazeCLIP framework
In order to apply the CLIP model to the gaze estimation, we design GazeCLIP as shown in Fig. 2. Firstly, we use CLIP pre-trained image encoder and text encoder to obtain their respective embeddings. Next, in order to effectively fuse the features of the two modalities instead of calculating similarity as in the CLIP model, we designed a cross attention mechanism to extract refined image embeddings that respond to text, and added them to the original image embeddings to enhance image representation. Finally, the enhanced image representation will be used for regression prediction to get gaze.
@@
Text Encoder. Due to the fact that the label in the gaze estimation is a 2-dimensional array corresponding to the pitch and yaw angles in the 3D gaze estimation, it is not appropriate to construct the prompt. We predefine the text prompt as “A photo of a face gazing [class]” to produce a universal phrase for images in gaze estimation dataset, and [class] denotes a group of primary directions (front, up, down, left, right). To guarantee that the prompt aligns with the language prior knowledge of the CLIP model, we employ the zero-shot CLIP model to allocate a direction to each image and the prompt composed of this direction has the highest cosine similarity to the current image. Meanwhile, we freeze the text encoder provided by CLIP to obtain well learned language prior knowledge. By tokenize and text encoder, here can obtain a text embedding T which is the same dimension with image embedding. This step can be defined as:
@@ 4.1 Datasets
MPIIFaceGaze. MPIIGaze [22] is the most commonly considered benchmark dataset for unconstrained 3D gaze estimation. And the MPIIFaceGaze is an extension of the MPIIGaze dataset, which provides 45K face images instead of eyes from 15 subjects captured by laptop camera during their daily life over several months. As a result, it contains images with diverse backgrounds, captured at varying times and under different lighting conditions. Leave-one-subject-out cross-validation is followed as the previous works do [11][17][23][25] which means leave one subject’s data as the test set at a time.
@@
RT-Gene. The RT-Gene dataset contains 122531 images of 15 subjects with wearable eye-tracking glasses which used for get gaze label. With the help of a series of wearable devices, the RT-gene dataset including a farther distance between the camera and the person, a greater variation in head poses and gazes compared to previous in-the-wild datasets. Therefore, this is a more challenging dataset. We follow the 3-fold evaluation protocol provided by RT-Gene dataset and divide the 15 subjects into three groups.
@@ 4.3 Implementation Details
We implement the experiments with PyTorch. All experiments are conducted on Nvidia A5000 GPU. The input images are 224×224×3 face images after normalizing. The backbone in our model is ViT-B/32 for image while Transformer[37] for text, they all generate features of the same dimension size in CLIP pre-trained model. And the parameter of text encoder is frozen during training. We train the GazeCLIP on three datasets with 128 batch size and 50 epochs. To optimize the model, we apply Adam optimizer [38] with an initial learning rate of 1e-5. And L1 Loss is applied for MPIIFaceGaze and EyeDiap, while L2 loss for RT-Gene. In order to adjust the learning rate during training process, we adopt the scheduler of MultiStepLR strategy. The learning rate will decay after the fifth epoch and 45th epoch, with a decay rate of 0.1.
@@
Improvement - $12\%$ $5\%$ $11\%$
Figure 3: Left: Performance on MPIIFaceGaze dataset. Middle: Performance on RT-Gene dataset. Right: Performance on EyeDiap dataset.
@@ 5.1 Comparison with the State-of-the-Arts
We first conduct an experiment to compare GazeCLIP with other state-of-the-art gaze estimation methods which are FullFace(spatial weights CNN)[17], Dilated-Net(a three streams network with dilated convolution)[24], RT-Gene( a three streams network with eyes and head pose)[12] and CA-Net(a coarse to fine adaptive network) [11]. Since RT-Gene report model ensemble is effective, here we present two results about RT-Gene (single model and 4 ensemble). Currently, the best performances in three datasets are achieved by CA-Net and RT-Gene (4 ensemble). Tab. 1 shows that GazeCLIP can achieve state-of-the-art results in all three datasets. Our method has improved by 0.5°($12\%$), 0.4°($5\%$), and 0.6°($11\%$) compared to the previous best results on three datasets, respectively. The excellent performance on three datasets demonstrates that our proposed GazeCLIP is a advantageous network architecture for gaze estimation.
@@ 5.2 Ablation Study
In order to further validate the effectiveness of the modules in the framework, we conducted three sets of ablation experiments. The following experiments are conducted on the MPIIFaceGaze[16] dataset, the most commonly used dataset in gaze estimation.
@@ 5.2.1 Ablation Study for language knowledge
We first analysis the effectiveness of language knowledge. We verify this through two strategies. Firstly, we maintain the existing network structure and change the pre-defined prompt to “ ” as text input; The second method is to remove the branch of language knowledge from the network and only use the image encoder for regression which makes the network no longer contain any text features. The results are shown in Tab. 2. If we only using the original CLIP image encoder with fine-tune like previous research architecture that only contain image features, however, reports unsatisfactory performance. Because more complex backbone often leads to more severe over-fitting problem in gaze estimation. When we keep the original network structure unchanged and replace the defined prompt with “ ”, the performance is significantly improved with 0.6°(4.4° → 3.8°). And taking a pre-defined prompt, “A photo of a face gazing [class]”, for a general language expression of the gaze estimation images as the input text achieves the best result. These results demonstrate the necessity of text features and the effectiveness of a appropriate prompt for text input in GazeCLIP.
@@ 5.2.2 Ablation Study for fixing different encoders
We further explore the effect of freezing image encoder and text encoder. As listed in Tab. 3, if we directly use the original CLIP model with a cross-attention mechanism and a regression head(i.e., the parameters of the image encoder and text encoder will not be optimized), we get the angular error of 8.9° which is not satisfactory. While we unfreeze the text encoder, it does not bring significant change and even worse(8.9° → 9.0°). However, when we choose to fine-tune the image encoder, here is a significant improvement with a drop form 8.9° to 3.8°. And if we try to fine-tune the text encoder as well, the angular error increased from 3.6° to 3.8°. Whether or not the image encoder is fixed, the fixed text encoder setting always performs better. We consider that the pre-trained CLIP model contains rich language priors that can encode natural languages well, and fine-tuning will break this. In addition, due to the data for CLIP pre-training come from a wide range of internet images, some special images such as in gaze estimation dataset may only account for a small part. As reported in [18], the zero-shot CLIP model can get a relative high accuracy in most image classification datasets but a simple dataset, MNIST[39]. Since the dataset for CLIP pre-training hardly contains handwritten digital images like the images in MNIST. Therefore, we need fine-tune the image encoder to cope with some particular tasks.
@@ 5.2.3 Ablation Study for feature fusion method
Then we study the influence of using different feature fusion methods during the training phase. In addition to using the cross attention mechanism we mentioned earlier, we also attempt two common methods: concat and add. Specifically, after obtaining image embedding and text embedding from CLIP encoders separately, we directly concatenate or add these two features. Since they have the same dimension, feature concat will double the feature dimension thus we need to modify the input dimension of the first linear layer in regression head, and feature add will not change the feature dimension, nor will cross attention mechanism. Results are shown in Tab. 4. Whether the feature fusion method is changed to concat or add, the angular error is 4.4°. Compared to the cross attention mechanism we use, it has increased by 0.8°. The results show that the cross attention mechanism is an effective way to establish an interaction between the language knowledge and image knowledge learned from the CLIP model.
@@ 5.3 Visual results
We display some visual images of our method in Fig. 5. These images indicate that our method can perform well in various situations, including subject gender, facial appearance, lighting conditions, whether glasses are worn , etc. We also presented some cases with poor results. As shown in Fig. 6, these images all have some special characteristics, such as partially closed eyes, dim facial lighting, blurred eye areas, etc. Due to not additional using eye images or paying more attention to the eye area, our method did not achieve satisfactory results on these images and we leave this problem as future work.
Moreover, as shown in the sixth sub-figure in Fig. 6 and Fig. 4, when take the 14th subject in the MPIIFaceGaze dataset as the test set, its experimental results show the highest angular error in 15-fold cross validation even if the appearance and lighting of the images are not very different from others. We speculate that the possible reason is that the images of the 14th subject always uses one hand to hold her face, which is almost non-existent in other subjects.
@@ 6 Conclusion
In this paper, we propose a novel network based on the vision-language pre-trained model (CLIP) to get accurate gaze estimation, named as GazeCLIP. In order to apply language knowledge to gaze estimation task, we uniformly establish corresponding language expressions for all images. And a key point is how to establish the connection between the two features. To make language knowledge effective, we designed a cross attention mechanism to enhance image representation. Experiments show the GazeCLIP can achieve state-of-the-art results in three well-known gaze estimation datasets ($12\%$, $5\%$, $11\%$improvement, respectively). We further conduct experiments to assess the advantages of different modules in our method. We hope this paper will provide a new insight for future research of gaze estimation with visual-language knowledge.

### v3

@@ 3.2 The GazeCLIP framework
In order to apply the CLIP model to the gaze estimation, we propose GazeCLIP, as shown in Fig. 2. Initially, we use the CLIP pre-trained image encoder and text encoder to obtain their respective embeddings. Subsequently, instead of simply computing similarity as in the CLIP model, we employ a cross-attention mechanism to effectively fuse the features of the two modalities. This mechanism extracts refined image embeddings that respond to text and combines them with the original image embeddings to enhance the image representation. Finally, enhanced image representation is used for regression prediction to estimate gaze directions.
@@
Text Encoder. Due to the fact that the label in the gaze estimation is a 2-dimensional array corresponding to the pitch and yaw angles in the 3D gaze estimation, it is not appropriate to construct the prompt. We pre-define the text prompt as ’A photo of a face gazing [class]’ to produce a universal phrase for images in the gaze estimation dataset, and [class] denotes a group of primary directions (front, up, down, left, right). To guarantee that the prompt aligns with the language prior knowledge of the CLIP model, we employ the zero-shot CLIP model to allocate a direction to each image and the prompt composed of this direction has the highest cosine similarity to the current image. Meanwhile, we freeze the text encoder provided by CLIP to obtain well-learned language prior knowledge. By tokenization and text encoder, here we can obtain a text embedding T which is the same dimension as the image embedding. This step can be defined as:
@@ 4.1 Datasets
MPIIFaceGaze. MPIIGaze [57] is the most commonly considered benchmark dataset for unconstrained 3D gaze estimation. And the MPIIFaceGaze is an extension of the MPIIGaze dataset, which provides 45K face images instead of eyes from 15 subjects captured by laptop camera during their daily life over several months. As a result, it contains images with diverse backgrounds, captured at varying times and under different lighting conditions. Leave-one-subject-out cross-validation is followed as the previous works do [8, 59, 7, 46] which means leaving one subject’s data as the test set at a time.
@@
RT-Gene. The RT-Gene dataset contains 122531 images of 15 subjects. With the help of a series of wearable devices, the RT-gene dataset including a farther distance between the camera and the person, a greater variation in head poses and gazes compared to previous in-the-wild datasets. Therefore, this is a more challenging dataset. We follow the 3-fold evaluation protocol provided by RT-Gene dataset and divide the 15 subjects into three groups.
@@ 4.3 Implementation Details
We implement the experiments with PyTorch. All experiments are conducted on Nvidia A5000 GPU. The input images are 224×224×3 face images after normalizing. The backbone in our model is ViT-B/32 for image while Transformer[45] for text, they all generate features of the same dimension size in CLIP pre-trained model. And the parameter of text encoder is frozen during training. We train GazeCLIP on three datasets with 128 batch sizes and 50 epochs. To optimize the model, we apply Adam optimizer [22] with an initial learning rate of 1e-5. And L1-Loss is applied for MPIIFaceGaze and EyeDiap, while L2-loss for RT-Gene. In order to adjust the learning rate during training process, we adopt the scheduler of MultiStepLR strategy. The learning rate will decay after the fifth epoch and 45th epoch, with a decay rate of 0.1.
@@
Improvement - $12\%$ $1\%$ $11\%$
@@ 5.1 Comparison with the State-of-the-Arts
We first conduct an experiment to compare GazeCLIP with other state-of-the-art gaze estimation methods. Since RT-Gene report model ensemble is effective, here we present two results about RT-Gene (single model and 4 ensemble). Currently, the best performance in three datasets is achieved by CA-Net and AGE-Net. Tab. 1 shows that GazeCLIP can achieve state-of-the-art results in all three datasets. Our method has improved by 0.5°($12\%$), 0.1°($1\%$), and 0.6°($11\%$) compared to the previous best results on three datasets, respectively. The excellent performance on three datasets demonstrates that our proposed GazeCLIP is an advantageous network architecture for gaze estimation.
@@ 5.2 Ablation Study
In order to further validate the effectiveness of the modules in our framework, we conduct three sets of ablation experiments. The following experiments are carried out on the MPIIFaceGaze[58] dataset, the most widely used dataset in gaze estimation.
@@ 5.2.1 Ablation Study for language knowledge
We first analyze the effectiveness of language knowledge through two strategies. Firstly, we maintain the existing network structure and change the pre-defined prompt to empty strings as text input; Secondly, we remove the branch of language knowledge from the network and only use the image encoder for regression, eliminating any text features from the network. This allows us to verify whether relying solely on pre-trained image encoders can achieve satisfactory performance. The results are summarized in Tab. 2. However, if we only keep the original CLIP image encoder as in previous research architectures that contain only visual features, reports unsatisfactory performance. Since a more complex backbone often leads to more severe over-fitting problem in gaze estimation when the dataset size is relatively small to the pre-trained dataset in CLIP. Interestingly, when we keep the original network structure unchanged and replace the defined prompt with empty strings, the performance is significantly improved by 0.6 ° (4.4 ° 3.8 °). And taking a pre-defined prompt, “A photo of a face gazing [class]”, for a general language expression of the gaze estimation images with rough directions as the input text achieves the best result. These results demonstrate the significance for text features and the effectiveness of an appropriate design of prompts in our framework.
@@ 5.2.2 Ablation Study for fixing different encoders
We further explore the effect of freezing image encoder and text encoder. As listed in Tab. 3, if we directly use the original CLIP model with a cross-attention mechanism and a regression head(i.e., the parameters of the image encoder and text encoder will not be optimized), we get the angular error of 8.9°. While we unfreeze the text encoder, it does not bring significant change and is even worse(8.9° → 9.0°). However, when we choose to fine-tune the image encoder, here is a significant improvement with a drop from 8.9° to 3.8°. And if we try to fine-tune the text encoder as well, the angular error increases from 3.6 ° to 3.8 °. Whether or not the image encoder is fixed, the fixed text encoder setting always performs better. We consider that the pre-trained CLIP model contains rich language priors that can encode natural languages well, and fine-tuning will break this. Furthermore, since the data for CLIP pre-training is drawn from a diverse range of internet images, certain characteristics present in gaze estimation datasets may be underrepresented. As reported in [35], the zero-shot CLIP model can get a relative high accuracy in most image classification datasets but a simple dataset, MNIST[25], since the dataset for CLIP pre-training hardly contains handwritten digital images like the images in MNIST. Therefore, we need fine-tune the image encoder to cope with some particular tasks.
@@ 5.2.3 Ablation Study for feature fusion method
Next, we study the influence of using different feature fusion methods during the training phase. In addition to using the cross-attention mechanism mentioned earlier, we also explore two common methods: concat and add. Specifically, after obtaining image embedding and text embedding from CLIP encoders separately, we directly concatenate or add these two features. Since they have the same dimension, feature concat will double the feature dimension. Then we need to modify the input dimension of the first linear layer in regression head, and feature add will not change the feature dimension, nor will cross-attention mechanism. The results are shown in Tab. 4. Whether the feature fusion method is changed to concat or add, the angular error is 4.4°. Compared to the cross-attention mechanism we use, it has increased by 0.8°. The findings demonstrate that the cross-attention mechanism serves as an effective method to establish an interaction between language knowledge and image knowledge acquired from the CLIP model.
@@ 5.3 Visual results
We display some visual images of our method in Fig. 4. These images indicate that our method can perform well in various situations, including subject gender, facial appearance, lighting conditions, whether glasses are worn , etc. We also presented some cases with poor results. As shown in Fig. 5, these images all have some special characteristics, such as partially closed eyes, dim facial lighting, blurred eye areas, etc. We attribute the unsatisfactory performance in these instances to our method’s lack of additional utilization of eye images or focused attention on the eye area. Addressing this limitation remains an area for future investigation.
Moreover, as shown in the sixth subfigure in Fig. 5 and Fig. 3, when taking the 14th subject in the MPIIFaceGaze dataset as the test set, its experimental results show the highest angular error in 15-fold cross-validation even if the appearance and lighting of the images are not very different from others. We posit that a potential explanation lies in the consistent depiction of the 14th subject holding her face with one hand, an occurrence largely absent in images of other subjects.
@@ 6 Conclusion
In this paper, we propose a novel network, GazeCLIP, based on the vision-language pre-trained model (CLIP) to get accurate gaze estimation. We leverage language knowledge by establishing uniform language expressions for all images, aiming to enhance the effectiveness of gaze estimation. A crucial aspect of our approach is establishing the connection between visual and textual features. To achieve this, we design a cross-attention mechanism to enhance image representation. Experimental results demonstrate that GazeCLIP achieves state-of-the-art performance across three well-known gaze estimation datasets, with improvements of ($12\%$, $1\%$, $11\%$ respectively).Furthermore, we conduct experiments to assess the advantages of different modules in our method. We believe this paper provides valuable insights for future research in gaze estimation leveraging visual-language knowledge.